OpenStudy (anonymous):

Ok, so the answer book doesn't have my answer: 3/4 x + 5/6 = 5x - 125/3. I get x = 15?

OpenStudy (amistre64):

does x=15 work?

OpenStudy (amistre64):

3/4 x + 5/6 = 5x - 125/3 ; multiply by 3 9/4 x + 15/6 = 15x - 125 ; since 15 and 6 have a common factor of 3, simplify 9/4 x + 5/2 = 15x - 125 ; multiply by 2 9(2)/4 x + 5 = 30x - 250 ; simplify the 2/4 9/2 x + 5 = 30x - 250 ; multiply by 2 again 9x + 10 = 60x - 500 ; solve for x 510 = 51x x = 10

OpenStudy (anonymous):

I almost had it!!! aaahhhhh.... Thanks so much...

OpenStudy (amistre64):

youre welcome :) there are of course other ways to approach this, but its all about not messing up lol